Suppose that the budget constraint is given as: PxX + PyY = M and the formulation of a utility function is given as:

χαΥβ U(X, Y)-2 log C + _ with 0 < α, β < 1 and constants C, δ > 0.

Answer for following questions and show all your calculation/ proof.

a. Derive the formula of income-consumption curve and draw its graph.

b. Derive the demand function of good X and Y respectively as functions of income and price of good X and Y.

c. Calculate the amount of income spent for good X and Y respectively.
Answer #1

Given, budget line: PxX + PyY = M
And utility function U(X,Y) = 2logC + rac{X^{alpha } + Y^{eta }}{delta } where C and delta are constants, and 0<alpha,beta<1

The budget line has a slope of Px

The Lagrangean approach transforms a constrained optimization problem into an unconstrained problem of choosing x, y, and the Lagrange multiplier,

to maximize L = u(x; y) + lambda[M - PxX - PyY]

By differentiating L with respect to x, y, and lambda, and setting the derivatives equal to zero, the resulting first order conditions
are:

rac{partial u}{partial x} - lambda P_{x} = 0

rac{partial u}{partial y} - lambda P_{y} = 0

M - P_{x}x - P_{y}y = 0
using the given utility function in the formulae above,
α一 - lambda P_{x}X= 0

BY8-1- lambda P_{y}Y = 0

M - P_{x}x - P_{y}y = 0

On solving, we get the required values of X and Y which form tangent to the utility curve and lie on the budget line. The income consumption curve for Px= 1 and Py=1 and M values of 4,5... looks like:

4 The Income-Consumption Curve

The graph includes the demand functions as well as the ICC curves. The quantities X and Y can be calculated using the method given above, after setting lambda = X = Y.